Simplify next expression t3-3t2s-s3+3ts2 and calculate if t=14, s=3


1
Expert's answer
2020-08-16T20:16:07-0400

It is given that we have to find expression value at

"t^3-3t^2s-s^3+3ts^2\\
when \,t=14 \, and \, s=3"


"t^3-s^3-3t^2s+3ts^2\\
=(t^3-s^3)-3ts(t-s)\\
=(t-s)(t^2+s^2+ts)-3ts(t-s)\\
=(t-s)(t^2+s^2-2ts)\\
(t-s)(t-s)^2\\
=(t-s)^3"

Now putting the value of t and s we get:

"(14-3)^3\\
=(11)^3\\
=1331"